OpenStudy (anonymous):

The manager of a baseball team has 15 players to choose from for his 9-person batting order. How many different ways can he arrange the players in the lineup? 5,005 362,880 3,603,600 1,816,214,400

OpenStudy (anonymous):

\[15\times 14\times 13\times 12\times 11\times 10\times 9\times 8\times 7\]

OpenStudy (anonymous):

1,816,214,400